# MUGen: Multi-modal Music Understanding and Generation with the Power of Large Language Models

### v4

@@ Abstract
The current landscape of research leveraging large language models (LLMs) is experiencing a surge. Many works harness the powerful reasoning capabilities of these models to comprehend various modalities, such as text, speech, images, videos, etc. They also utilize LLMs to understand human intention and generate desired outputs like images, videos and music. However, research that combines both understanding and generation using LLMs is still limited and in its nascent stage. To address this gap, we introduce a Multi-modal Music Understanding and Generation (MUGen) framework that integrates LLM’s abilities to comprehend and generate music for different modalities. The MUGen framework is purpose-built to unlock creative potential from diverse sources of inspiration, encompassing music, image and video through the use of pretrained MERT, ViT, and ViViT models, respectively. To enable music generation, we explore the use of AudioLDM 2 and MusicGen. Bridging multi-modal understanding and music generation is accomplished through the integration of the LLaMA 2 model. Furthermore, we make use of the MU-LLaMA model to generate extensive datasets that support text/image/video-to-music generation, facilitating the training of our MUGen framework. We conduct a thorough evaluation of our proposed framework. The experimental results demonstrate that our model achieves or surpasses the performance of the current state-of-the-art models.
![[Uncaptioned image]](extracted/5449010/assets/Cover.png)
@@ 1 Introduction
In this work, we propose the Multi-modal Music Understanding and Generation (MUGen) framework capable of music understanding and drawing inspiration from multi-modal inputs to generate music. Our framework employs multiple modal encoders to represent image, video and music inputs, respectively. In line with the conventions of the computer vision field, we select ViT [15] and ViViT [4] as the encoders for the image and video modalities. The MERT model [41], which has exhibited outstanding performance in downstream music tagging tasks in the MU-LLaMA work [47], is chosen as our music encoder. The feature representations obtained from the input encoders of different modalities are then fed into their respective understanding adaptors. The LLaMA 2 model [65] comprehends these modality signals and input intentions to carry out downstream tasks. For the music generation task, we explore and compare two music decoders, which are AudioLDM 2 [46] and MusicGen [9], while music understanding is directly addressed by the LLaMA 2 model.
@@
We introduce the MUGen framework, an advancement capable of simultaneously encompassing music understanding and multi-modal music generation tasks, aiming to assist users in music related artistic creation.
@@
We propose a systematic approach for generating large multi-modal music oriented instruction datasets for the training of our MUGen model.
@@ 2 Related Works
Research in music generation has made significant strides in recent years, particularly with the rise of Transformer [67] and diffusion [29] models. Some notable works representing the utilization of text as input for guiding music generation include MusicLM [1], MusicGen [9], Moûsai [57], Noise2music [32], and AudioLDM 2 [46]. Among these, AudioLDM 2 and MusicGen stand out for their exceptional music generation abilities. AudioLDM 2 is a general-purpose audio generation model based on the diffusion process, while MusicGen consists of a single autoregressive Transformer decoder. In the proposed MUGen framework, we explore and compare these two music decoders. There have been a few works in investigating music generation guided by alternative modalities, such as Vis2Mus [81] for generating music from images, and CMT [11] and V-MusPod [87] for generating music from videos. However, these approaches are limited to a single modality as the instruction condition. In contrast, this paper presents a novel approach that integrates multiple modalities, including text, image, and video, leveraging the LLM framework to achieve multi-modal music generation.
@@
MLLMs have emerged as a prominent research topic, with researchers tirelessly equipping these models with various functionalities. For example, Macaw-LLM [49] integrates text, images, videos, and audio into a unified framework, enabling multi-modal comprehension. DreamLLM [14] and InternLM-XComposer [80] generate text and images in an interleaved manner, enhancing the multi-modal conversational abilities between users and LLMs. For instance, InternLM-XComposer can compose an article with accompanying illustrations. SEED-LLaMA [21] leverages LLaMA [65] and diffusion models for image understanding and question answering, as well as image generation. The NExT-GPT model [71], which is most relevant to our proposed framework, facilitates interleaved conversations involving text, images, videos, and audio. However, its limited music training data restricts its musical capabilities. In this paper, we introduce the MUGen framework, which enables music understanding and multi-modal music generation guided by multiple modalities. It also allows for modifying input music based on music prompts and text. We believe that our work will make a valuable contribution to the community of AI-assisted artistic creation.
3 MUGen Model Architecture & Training
The architecture of the MUGen model is illustrated in Figure 2. In this section, we provide an in-depth description of the MUGen architecture and elaborate on the training methods employed during the development of this model.
@@ 3.1 Multi-modal Feature Encoders
In order to accomplish multi-modal music understanding and generation, the MUGen model needs to be able to process multi-modal inputs. To achieve this, it incorporates frozen pre-trained encoders capable of comprehending various modalities, including music, images, and videos. These encoders extract relevant information to address tasks associated with music question answering and music generation within the MUGen framework. Specifically, the following encoders are employed for processing involved modalities:
@@ 3.4 Music Understanding and Generation
Inspired by models such as NExT-GPT [71], the MUGen model incorporates specialized audio tokens of the form , where , to distinguish between music question answering and generation tasks. The number of audio tokens is a hyper-parameter that determines the dimension of the input to the music output Transformer, also known as the output projector, during music generation. In the training phase, instruction sets containing music as the output are adjusted by appending such audio tokens to the end of the output. In the inference phase, the MUGen model will generate audio tokens only if the instruction prompt requires a music output. Through this method, the MUGen model exhibits the capability to produce text-only outputs for music question answering and music outputs for music generation within the same framework.
@@ 3.5 Training Method
where denotes target output tokens, denotes the target embeddings for AudioLDM 2/MusicGen, represents the output from MUGen’s LLaMA 2 model, represents the output from MUGen’s output projection layer, and is the cross entropy (CE) loss. The loss function used by the MUGen model is either CE alone or a combination of CE and mean squared error (MSE). During training, if the task involves only text token generation, the model is guided by the CE loss alone. If the task involves music generation, both CE and MSE are used together, with MSE calculated between the conditioning embedding generated by the output projector and the target music caption’s text encoding from the music generation model. This dual-loss strategy ensures that the MUGen model is adept at both text token generation and generating embeddings for conditioning the downstream music generation models (AudioLDM 2 or MusicGen).
@@ 4 Music Oriented Instruction Dataset
To train our MUGen model, the availability of multi-modal instruction datasets that support any-to-music generation is imperative. Moreover, extensive datasets that include pairs of different modalities, such as text-image pairs, are essential for modality alignment training. We make use of established datasets like Alpaca [62] with general knowledge for instruction following and COCO [45] for aligning the image encoder. In addition to utilizing existing resources, we also collect our own dataset. We adopt an automated approach to overcome the labor-intensive and time-consuming nature of manual data collection. Specifically, inspired by previous works [47, 25], the MU-LLaMA [47] and MPT-7B [63] models are utilized to generate required datasets. In the following subsections, we provide a comprehensive overview of the methodologies employed in crafting the datasets used for training the MUGen model.
@@ 4.4 MUVideo Dataset
Similar to MUImage, we also create the MUVideo dataset to enable our proposed MUGen framework to accomplish the task of video-to-music generation. The MUVideo dataset is curated by gathering music samples from the AudioSet as well with their corresponding videos. To construct this dataset, the following steps are designed:
@@ 5 Model Evaluation
Given various capabilities demonstrated by our MUGen model, such as music understanding and music generation from multi-modal inputs, we conduct a comprehensive evaluation of the model in this section, assessing its performance across different subtasks. We also present a comparative analysis with other pertinent models. One such model demonstrating the capability of any-to-any generation is NExT-GPT[71]; however, since the checkpoint released by the authors can not function as expected and has issues generating desired outputs, a direct comparison with NExT-GPT for large-scale evaluation is currently unfeasible. During the evaluation, we set the hyper-parameters of the MUGen model as follows: temperature, top_p and max target length. We have also made sure that all models leveraging LLMs, such as LLaMA-Adapter [18] and SALMONN [60], use the same hyper-parameters for evaluation to ensure a fair comparison.
@@ 5.1 Music Understanding
To evaluate the music understanding capabilities of the MUGen model, we employ the MTG-eval-QA subset of the MusicQA dataset proposed by Liu et al. (2023) [47] as our evaluation set, consisting of 4,500 music question-answer pairs. The SOTA models for comparison include LTU [25], LLaMA-Adapter [18], SALMONN [60] and MU-LLaMA [47]. Among these models, MU-LLaMA stands out as the only one that was specifically trained on music-related datasets. The evaluation metrics used for music understanding follow the work of [47], containing BLEU (B-U) [56], METEOR (M-R) [6], ROUGE (R-L) [44], and BERT-Score (BERT-S) [83].
@@
MUGen | 0.308 | 0.393 | 0.476 | 0.902 |
It is evident from the Table 1 that the LTU, LLaMA Adapter, and SALMONN models are lagging behind MU-LLaMA and our MUGen model in the four metrics compared, since the former three models have not been specifically trained on music-related datasets. Both the MU-LLaMA and our MUGen models have been trained on the MusicQA dataset [47], demonstrating impressive music understanding capabilities. It is worth noting that our model still outperforms the MU-LLaMA model. One possible reason is that, in addition to the MusicQA dataset, we also have the MUCaps dataset for text-music alignment training, which contributes to enhancing our model’s music understanding abilities. However, since the MUCaps dataset is generated by the MU-LLaMA model, there may be limitations on further improving our model’s music understanding capabilities. This could be one of the factors preventing our model from significantly surpassing MU-LLaMA, which implies that in our future work, we may need to explore other music datasets to make further improvements.
@@ 5.2 Text to Music Generation
For text-to-music generation, we use the evaluation set from the MUCaps dataset. This set comprises 5,000 text-music pairs. SOTA models selected for comparison include CoDi [61], AudioLDM 2 [46], and MusicGen [9]. Among these models, MusicGen is the sole one explicitly trained for music generation, leading us to anticipate its superior performance compared to other models. In this and the following experiments, we evaluate the MUGen model with both AudioLDM 2 (MUGen v1) [46] and MusicGen (MUGen v2) [9] decoders. Following MusicGen, we use the Fréchet Audio Distance (FAD) [37], Kullback-Leibler divergence (KL), and CLAP score [72] as the evaluation metrics.
Model | FAD | KL | CLAP |
@@
MUGen v1 | 11.143 | 3.982 | 0.282 |
MUGen v2 | 10.498 | 3.769 | 0.313 |
From Table 2, we can see that CoDi performs the worst of all the models for the task of text-to-music generation. This is likely due to the fact that it has not been trained on music data. As mentioned above, our model incorporates AudioLDM 2 and MusicGen as music decoders, so in Table 2, MUGen v1 is compared to AudioLDM 2, and MUGen v2 is compared to MusicGen. It can be observed that our MUGen model performs better when given AudioLDM 2 or MusicGen as the music decoder compared to using them alone. Particularly, the generated music is more relevant to the input instructions, as indicated by the improvement in CLAP score. This can be attributed to the use of LLMs, which makes it easier for the model to understand the input instructions and use them to guide music generation.
@@ 5.3 Prompt Based Music Editing
MUGen is one of the few existing models that support music editing using natural language. In contrast, models like AUDIT [68] and InstructME [27] require specific prompt words such as “Add”, “Drop”, “Remove” and others to edit music. Loop Copilot [85] is another model that supports music editing using natural language. However, since the authors have not made their model open-source, we are unable to compare it. Both AUDIT and InstructME also lack open-sourced models, but InstructME provides a few samples that we can utilize for comparison. Following AUDIT, in addition to using FAD and KL for evaluation, we introduce another metric called log spectral distance (LSD) to assess the music editing subtask.
Model | FAD | KL | LSD |
@@
MUGen v1 | 2.223 | 5.654 | 0.790 |
MUGen v2 | 2.191 | 5.118 | 0.735 |
Table 3 illustrates the superior performance of our MUGen model compared to the AUDIT and InstructME models. This achievement can be attributed to the utilization of the LLaMA 2 model for comprehending prompts in the editing task. This allows the model to effectively edit music based on natural language prompts. Furthermore, the use of the MERT Encoder enhances the model’s capability to better understand the source music, consequently improving its effectiveness during the editing process.
Model | FAD | KL | IB Rank |
@@
MUGen v1 | 7.326 | 6.014 | 0.688 |
MUGen v2 | 6.968 | 5.878 | 0.819 |
Model | FAD | KL | IB Rank |
@@ 5.4 Multi-modal Music Generation
Multi-modal music generation from images/videos is a crucial ability of our MUGen model. CoDi [61] is an any-to-any generation model, thus it is involved for both image-to-music (I2M) and video-to-music (V2M) generation experiments (Tables 4 and 5), while CMT [11] is for video-to-music generation. The evaluation sets for I2M and V2M consist of 2,500 pairs of image-music and video-music, respectively. In these two sets of experiments, apart from FAD and KL, we introduce a new evaluation metric called ImageBind Ranking (IB Rank) [23] to assess the alignment between the image/video modality and the generated music. Specifically, we use the ImageBind model to obtain embeddings for the images/videos and the generated music, and then calculate their similarity scores for ranking purposes.

### v5

# M2UGen: Multi-modal Music Understanding and Generation with the Power of Large Language Models

@@ Abstract
The current landscape of research leveraging large language models (LLMs) is experiencing a surge. Many works harness the powerful reasoning capabilities of these models to comprehend various modalities, such as text, speech, images, videos, etc. They also utilize LLMs to understand human intention and generate desired outputs like images, videos and music. However, research that combines both understanding and generation using LLMs is still limited and in its nascent stage. To address this gap, we introduce a Multi-modal Music Understanding and Generation (M2UGen) framework that integrates LLM’s abilities to comprehend and generate music for different modalities. The M2UGen framework is purpose-built to unlock creative potential from diverse sources of inspiration, encompassing music, image and video through the use of pretrained MERT, ViT, and ViViT models, respectively. To enable music generation, we explore the use of AudioLDM 2 and MusicGen. Bridging multi-modal understanding and music generation is accomplished through the integration of the LLaMA 2 model. Furthermore, we make use of the MU-LLaMA model to generate extensive datasets that support text/image/video-to-music generation, facilitating the training of our M2UGen framework. We conduct a thorough evaluation of our proposed framework. The experimental results demonstrate that our model achieves or surpasses the performance of the current state-of-the-art models.
Please refer to and cite the latest version by searching for the title: “MuMu-LLaMA: Multi-modal Music Understanding
and Generation via Large Language Models”.
@@ 1 Introduction
In this work, we propose the Multi-modal Music Understanding and Generation (M2UGen) framework capable of music understanding and drawing inspiration from multi-modal inputs to generate music. Our framework employs multiple modal encoders to represent image, video and music inputs, respectively. In line with the conventions of the computer vision field, we select ViT [15] and ViViT [4] as the encoders for the image and video modalities. The MERT model [41], which has exhibited outstanding performance in downstream music tagging tasks in the MU-LLaMA work [47], is chosen as our music encoder. The feature representations obtained from the input encoders of different modalities are then fed into their respective understanding adaptors. The LLaMA 2 model [65] comprehends these modality signals and input intentions to carry out downstream tasks. For the music generation task, we explore and compare two music decoders, which are AudioLDM 2 [46] and MusicGen [9], while music understanding is directly addressed by the LLaMA 2 model.
@@
We introduce the M2UGen framework, an advancement capable of simultaneously encompassing music understanding and multi-modal music generation tasks, aiming to assist users in music related artistic creation.
@@
We propose a systematic approach for generating large multi-modal music oriented instruction datasets for the training of our M2UGen model.
@@ 2 Related Works
Research in music generation has made significant strides in recent years, particularly with the rise of Transformer [67] and diffusion [29] models. Some notable works representing the utilization of text as input for guiding music generation include MusicLM [1], MusicGen [9], Moûsai [57], Noise2music [32], and AudioLDM 2 [46]. Among these, AudioLDM 2 and MusicGen stand out for their exceptional music generation abilities. AudioLDM 2 is a general-purpose audio generation model based on the diffusion process, while MusicGen consists of a single autoregressive Transformer decoder. In the proposed M2UGen framework, we explore and compare these two music decoders. There have been a few works in investigating music generation guided by alternative modalities, such as Vis2Mus [81] for generating music from images, and CMT [11] and V-MusPod [87] for generating music from videos. However, these approaches are limited to a single modality as the instruction condition. In contrast, this paper presents a novel approach that integrates multiple modalities, including text, image, and video, leveraging the LLM framework to achieve multi-modal music generation.
@@
MLLMs have emerged as a prominent research topic, with researchers tirelessly equipping these models with various functionalities. For example, Macaw-LLM [49] integrates text, images, videos, and audio into a unified framework, enabling multi-modal comprehension. DreamLLM [14] and InternLM-XComposer [80] generate text and images in an interleaved manner, enhancing the multi-modal conversational abilities between users and LLMs. For instance, InternLM-XComposer can compose an article with accompanying illustrations. SEED-LLaMA [21] leverages LLaMA [65] and diffusion models for image understanding and question answering, as well as image generation. The NExT-GPT model [71], which is most relevant to our proposed framework, facilitates interleaved conversations involving text, images, videos, and audio. However, its limited music training data restricts its musical capabilities. In this paper, we introduce the M2UGen framework, which enables music understanding and multi-modal music generation guided by multiple modalities. It also allows for modifying input music based on music prompts and text. We believe that our work will make a valuable contribution to the community of AI-assisted artistic creation.
3 M2UGen Model Architecture & Training
The architecture of the M2UGen model is illustrated in Figure 2. In this section, we provide an in-depth description of the M2UGen architecture and elaborate on the training methods employed during the development of this model.
@@ 3.1 Multi-modal Feature Encoders
In order to accomplish multi-modal music understanding and generation, the M2UGen model needs to be able to process multi-modal inputs. To achieve this, it incorporates frozen pre-trained encoders capable of comprehending various modalities, including music, images, and videos. These encoders extract relevant information to address tasks associated with music question answering and music generation within the M2UGen framework. Specifically, the following encoders are employed for processing involved modalities:
@@ 3.4 Music Understanding and Generation
Inspired by models such as NExT-GPT [71], the M2UGen model incorporates specialized audio tokens of the form , where , to distinguish between music question answering and generation tasks. The number of audio tokens is a hyper-parameter that determines the dimension of the input to the music output Transformer, also known as the output projector, during music generation. In the training phase, instruction sets containing music as the output are adjusted by appending such audio tokens to the end of the output. In the inference phase, the M2UGen model will generate audio tokens only if the instruction prompt requires a music output. Through this method, the M2UGen model exhibits the capability to produce text-only outputs for music question answering and music outputs for music generation within the same framework.
@@ 3.5 Training Method
where denotes target output tokens, denotes the target embeddings for AudioLDM 2/MusicGen, represents the output from M2UGen’s LLaMA 2 model, represents the output from M2UGen’s output projection layer, and is the cross entropy (CE) loss. The loss function used by the M2UGen model is either CE alone or a combination of CE and mean squared error (MSE). During training, if the task involves only text token generation, the model is guided by the CE loss alone. If the task involves music generation, both CE and MSE are used together, with MSE calculated between the conditioning embedding generated by the output projector and the target music caption’s text encoding from the music generation model. This dual-loss strategy ensures that the M2UGen model is adept at both text token generation and generating embeddings for conditioning the downstream music generation models (AudioLDM 2 or MusicGen).
@@ 4 Music Oriented Instruction Dataset
To train our M2UGen model, the availability of multi-modal instruction datasets that support any-to-music generation is imperative. Moreover, extensive datasets that include pairs of different modalities, such as text-image pairs, are essential for modality alignment training. We make use of established datasets like Alpaca [62] with general knowledge for instruction following and COCO [45] for aligning the image encoder. In addition to utilizing existing resources, we also collect our own dataset. We adopt an automated approach to overcome the labor-intensive and time-consuming nature of manual data collection. Specifically, inspired by previous works [47, 25], the MU-LLaMA [47] and MPT-7B [63] models are utilized to generate required datasets. In the following subsections, we provide a comprehensive overview of the methodologies employed in crafting the datasets used for training the M2UGen model.
@@ 4.4 MUVideo Dataset
Similar to MUImage, we also create the MUVideo dataset to enable our proposed M2UGen framework to accomplish the task of video-to-music generation. The MUVideo dataset is curated by gathering music samples from the AudioSet as well with their corresponding videos. To construct this dataset, the following steps are designed:
@@ 5 Model Evaluation
Given various capabilities demonstrated by our M2UGen model, such as music understanding and music generation from multi-modal inputs, we conduct a comprehensive evaluation of the model in this section, assessing its performance across different subtasks. We also present a comparative analysis with other pertinent models. One such model demonstrating the capability of any-to-any generation is NExT-GPT[71]; however, since the checkpoint released by the authors can not function as expected and has issues generating desired outputs, a direct comparison with NExT-GPT for large-scale evaluation is currently unfeasible. During the evaluation, we set the hyper-parameters of the M2UGen model as follows: temperature, top_p and max target length. We have also made sure that all models leveraging LLMs, such as LLaMA-Adapter [18] and SALMONN [60], use the same hyper-parameters for evaluation to ensure a fair comparison.
@@ 5.1 Music Understanding
To evaluate the music understanding capabilities of the M2UGen model, we employ the MTG-eval-QA subset of the MusicQA dataset proposed by Liu et al. (2023) [47] as our evaluation set, consisting of 4,500 music question-answer pairs. The SOTA models for comparison include LTU [25], LLaMA-Adapter [18], SALMONN [60] and MU-LLaMA [47]. Among these models, MU-LLaMA stands out as the only one that was specifically trained on music-related datasets. The evaluation metrics used for music understanding follow the work of [47], containing BLEU (B-U) [56], METEOR (M-R) [6], ROUGEL (R-L) [44], and BERT-Score (BERT-S) [83].
@@
| M2UGen | 0.308 | 0.393 | 0.476 | 0.902 |
It is evident from the Table 1 that the LTU, LLaMA Adapter, and SALMONN models are lagging behind MU-LLaMA and our M2UGen model in the four metrics compared, since the former three models have not been specifically trained on music-related datasets. Both the MU-LLaMA and our M2UGen models have been trained on the MusicQA dataset [47], demonstrating impressive music understanding capabilities. It is worth noting that our model still outperforms the MU-LLaMA model. One possible reason is that, in addition to the MusicQA dataset, we also have the MUCaps dataset for text-music alignment training, which contributes to enhancing our model’s music understanding abilities. However, since the MUCaps dataset is generated by the MU-LLaMA model, there may be limitations on further improving our model’s music understanding capabilities. This could be one of the factors preventing our model from significantly surpassing MU-LLaMA, which implies that in our future work, we may need to explore other music datasets to make further improvements.
@@ 5.2 Text to Music Generation
For text-to-music generation, we use the evaluation set from the MUCaps dataset. This set comprises 5,000 text-music pairs. SOTA models selected for comparison include CoDi [61], AudioLDM 2 [46], and MusicGen [9]. Among these models, MusicGen is the sole one explicitly trained for music generation, leading us to anticipate its superior performance compared to other models. In this and the following experiments, we evaluate the M2UGen model with both AudioLDM 2 (M2UGen v1) [46] and MusicGen (M2UGen v2) [9] decoders. Following MusicGen, we use the Fréchet Audio Distance (FAD) [37], Kullback-Leibler divergence (KL), and CLAP score [72] as the evaluation metrics.
| Model | FADvgg | KL | CLAPscore |
@@
| M2UGen v1 | 11.143 | 3.982 | 0.282 |
| M2UGen v2 | 10.498 | 3.769 | 0.313 |
From Table 2, we can see that CoDi performs the worst of all the models for the task of text-to-music generation. This is likely due to the fact that it has not been trained on music data. As mentioned above, our model incorporates AudioLDM 2 and MusicGen as music decoders, so in Table 2, M2UGen v1 is compared to AudioLDM 2, and M2UGen v2 is compared to MusicGen. It can be observed that our M2UGen model performs better when given AudioLDM 2 or MusicGen as the music decoder compared to using them alone. Particularly, the generated music is more relevant to the input instructions, as indicated by the improvement in CLAP score. This can be attributed to the use of LLMs, which makes it easier for the model to understand the input instructions and use them to guide music generation.
@@ 5.3 Prompt Based Music Editing
M2UGen is one of the few existing models that support music editing using natural language. In contrast, models like AUDIT [68] and InstructME [27] require specific prompt words such as “Add”, “Drop”, “Remove” and others to edit music. Loop Copilot [85] is another model that supports music editing using natural language. However, since the authors have not made their model open-source, we are unable to compare it. Both AUDIT and InstructME also lack open-sourced models, but InstructME provides a few samples that we can utilize for comparison. Following AUDIT, in addition to using FAD and KL for evaluation, we introduce another metric called log spectral distance (LSD) to assess the music editing subtask.
| Model | FADvgg | KL | LSD |
@@
| M2UGen v1 | 2.223 | 5.654 | 0.790 |
| M2UGen v2 | 2.191 | 5.118 | 0.735 |
Table 3 illustrates the superior performance of our M2UGen model compared to the AUDIT and InstructME models. This achievement can be attributed to the utilization of the LLaMA 2 model for comprehending prompts in the editing task. This allows the model to effectively edit music based on natural language prompts. Furthermore, the use of the MERT Encoder enhances the model’s capability to better understand the source music, consequently improving its effectiveness during the editing process.
| Model | FADvgg | KL | IB Rank |
@@
| M2UGen v1 | 7.326 | 6.014 | 0.688 |
| M2UGen v2 | 6.968 | 5.878 | 0.819 |
| Model | FADvgg | KL | IB Rank |
@@ 5.4 Multi-modal Music Generation
Multi-modal music generation from images/videos is a crucial ability of our M2UGen model. CoDi [61] is an any-to-any generation model, thus it is involved for both image-to-music (I2M) and video-to-music (V2M) generation experiments (Tables 4 and 5), while CMT [11] is for video-to-music generation. The evaluation sets for I2M and V2M consist of 2,500 pairs of image-music and video-music, respectively. In these two sets of experiments, apart from FAD and KL, we introduce a new evaluation metric called ImageBind Ranking (IB Rank) [23] to assess the alignment between the image/video modality and the generated music. Specifically, we use the ImageBind model to obtain embeddings for the images/videos and the generated music, and then calculate their similarity scores for ranking purposes.